OpenStudy (anonymous):

Solve: −4(x + 10) − 6 = −3(x − 2)

OpenStudy (goldphenoix):

Do you know how to simplify the right side?

OpenStudy (anonymous):

-3x=6

OpenStudy (uri):

I am telling youu i know tyhe ansurrr!!

OpenStudy (anonymous):

what is it?

OpenStudy (uri):

Solve the bracket first @shahbaaz786

OpenStudy (anonymous):

what do you mean the bracket?

OpenStudy (uri):

−4(x + 10) − 6 = −3(x − 2) -4x-40-6=-3x+6

OpenStudy (uri):

Now put all x values together and all numericals together

OpenStudy (anonymous):

x=52?

OpenStudy (anonymous):

-52?

OpenStudy (uri):

-4x-40-6=-3x+6 -4x+3x=6+40+6 -x=52 x=-52

OpenStudy (goldphenoix):

Your ss went up, urie. :) Congrantz.

OpenStudy (uri):

My head aches so bad right now!! I should go sleep.

OpenStudy (uri):

Oh because I'm smart...nowadays....only in holidays...my life saver went up too!! wait yesterday only my SS turned to 82?NO!

OpenStudy (goldphenoix):

Go to sleep. :)

OpenStudy (uri):

I think they will kick me out.